OpenStudy (anonymous):

how do u read this and what does it mean? for function y(x) it holds that at (x,y) = (1,y(1))

OpenStudy (freckles):

that means when x=1 , y=y(1)

OpenStudy (freckles):

for example pretend y=2x+1 y is a function of x so i will write y=y(x) y(x)=2x+1 when x=1 we have y=y(1) so replace x with 1 to find y(1) y(1)=2(1)+1 y(1)=2+1 y(1)=3 so this means when x=1 we have y=y(1)=3

OpenStudy (freckles):

you know you might have seen (x,y)=(x,f(x)) before

OpenStudy (freckles):

instead of calling y f we could also just call it y

OpenStudy (anonymous):

thnx man!